### v1

# spacing=nonfrench

Abstract
Anti-malware systems rely on sandboxes, hooks, and telemetry pipelines, including collection agents, serializers, and database backends, to monitor program and system behavior. We show that these data-handling components constitute an exploitable attack surface that can lead to denial-of-analysis (DoA) states without disabling sensors or requiring elevated privileges. As a result, we present Telemetry Complexity Attacks (TCAs), a new class of vulnerabilities that exploit fundamental mismatches between unbounded collection mechanisms and bounded processing capabilities. Our method recursively spawns child processes to generate specially crafted, deeply nested, and oversized telemetry that stresses serialization and storage boundaries, as well as visualization layers, for example, JSON/BSON depth and size limits. Depending on the product, this leads to truncated or missing behavioral reports, rejected database inserts, serializer recursion and size errors, and unresponsive dashboards. In all of these cases, malicious activity is normally executed; however, depending on the examined solution, it is not recorded and/or not presented to the analysts. Therefore, instead of evading sensors, we break the pipeline that stores the data captured by the sensors.
We evaluate our technique against twelve commercial and open-source malware analysis platforms and endpoint detection and response (EDR) solutions. Seven products fail in different stages of the telemetry pipeline; two vendors assigned CVE identifiers (CVE-2025-61301 and CVE-2025-61303), and others issued patches or configuration changes. We discuss root causes and propose mitigation strategies to prevent DoA attacks triggered by adversarial telemetry.
1 Introduction
The evolution of malware and defense tooling has led to an ongoing arms race, where each side continuously deploys new and increasingly sophisticated mechanisms to counter the other. On one side, modern malware analysis and endpoint protection tools rely on telemetry pipelines to capture program behavior. These pipelines combine process monitoring, dynamic sandboxes, API hooking, ETW/Sysmon instrumentation, event collectors, host-based collection agents, serialization layers, and storage backends. Their purpose is to observe execution and produce behavioral reports that analysts and automated classifiers can inspect. These steps are assumed to be reliable; once a process is observed, its behavior is expected to appear in the resulting analysis report.
In response, adversaries invest in anti-analysis techniques that detect virtualized or instrumented environments and suppress or alter execution accordingly. While this line of work is important, it primarily focuses on avoiding observation by evading or disabling the sensing substrate. In this regard, these approaches are inherently evasive, as the goal is not to overload the monitoring system, but to leave no meaningful telemetry in the first place. However, we identify a critical assumption in many anti-malware solutions: telemetry pipelines treat monitoring data as trusted input rather than a potential attack vector.
Therefore, our insight is that the telemetry pipeline itself constitutes an attack surface. As a result, we introduce Telemetry Complexity Attacks (TCAs), which exploit the semantic gap between unbounded behavioral observation and bounded data processing capabilities. Unlike evasion techniques that hide malicious activity, TCAs generate pathological but valid telemetry that overwhelms serialization, storage, and rendering components. This creates a ‘denial-of-analysis’ condition where malicious activity executes successfully but remains invisible to defenders.
More precisely, we examine attacks that target the telemetry pipeline itself to induce a denial-of-analysis condition, focusing on the data representation. Instead of avoiding observation (evasion) or modifying sensors (anti-hooking), we target a different part of the pipeline: the data structures used to represent telemetry. By crafting execution behavior that produces deeply nested or unusually voluminous telemetry, we push serialization libraries and storage engines into failure states. These include JSON and BSON encoders, serialization libraries, document databases, and UI rendering subsystems. When internal limits are reached, for example, maximum nesting depth, recursion limits, or document size caps, the telemetry pipeline collapses into a denial-of-analysis state, that is, analysis continues, the malicious activity executes, but the resulting report is incomplete, truncated, or entirely missing.
To provide the reader with some further insight, we consider a malware sample that recursively spawns child processes to a depth of 200. While the operating system handles this load, modern systems can support $\sim$32K processes, the telemetry pipeline must:
  1. 1.
    Serialize a process tree.
  2. 2.
    Store this in JSON/BSON with 100-level nesting limits.
  3. 3.
    Render it on dashboards with finite DOM constraints.
The mismatch between collection and processing capabilities creates exploitable failure modes at multiple stages of the pipeline.
Through a thorough and detailed experimental process, we demonstrate that this vulnerability is present in several real-world anti-malware products and services. In our experiments, the main observed impact includes:
  • missing or partial behavioral reports,
  • rejected database inserts due to size or depth limits,
  • serializer crashes caused by recursion or memory exhaustion, and
  • stalled or unresponsive dashboards.
Notably, these effects require no elevated privileges and do not tamper with hooks, sensors, or agents. They arise solely from adversarially structured telemetry that is valid, but pathological.
We validate our approach across commercial and open-source products. Several vendors confirmed the issue and deployed mitigations, and, at the time of writing, two vulnerabilities have received CVE identifiers.
1.1 Responsible disclosure
We disclosed the issues to the affected vendors and coordinated public disclosure in accordance with standard practices, providing at least 90 days for them to patch their solutions. At the time of writing, two issues have been assigned CVE IDs; namely CVE-2025-61301 and CVE-2025-61303, while other vendors have deployed configuration changes or patches. Proof-of-concept artifacts for the two CVEs are available on GitHub [^1] https://github.com/eGkritsis/CVE-2025-61301/ and https://github.com/eGkritsis/CVE-2025-61303/, respectively.
1.2 Contribution
The core contribution of this work is the introduction of a new attack to anti-malware solutions, TCAs. We formalize TCAs as distinct from resource exhaustion attacks, with unique failure modes in telemetry pipelines. Our approach is straightforward. Instead of trying to disable hooks or evade instrumentation, we manipulate the data that is monitored to construct telemetry records that push ingestion and storage components into failure. By abusing the container formats and serialization boundaries that anti-malware systems depend on, we can provoke faults at multiple points in the pipeline, e.g., during ingestion, transformation, or storage, with outcomes ranging from missing fields in reports to complete pipeline collapse. We tested this technique against several anti-malware solutions, leading to several patches in the corresponding solutions and the publication of two CVEs. Table 1 summarizes the results and the severity of the failures observed.
We developed a methodology that recursively spawns child processes to generate high-volume telemetry, stressing serialization and storage boundaries and causing analysis failure or denial-of-service in downstream components. We analyze the approach and provide a binary proof-of-concept able to reach dynamic analysis engines (sandbox or EDR telemetry pipelines) without being classified as malware by static detection. Once executed in the monitored environment, the binary stresses or corrupts the product’s data-handling pipeline (e.g., log generation, serialization, storage, indexing, or UI rendering) so that the resulting report is produced incompletely, corrupted, truncated, or the analysis UI is rendered unusable, depending on the software vendor and the underlying stack.
1.3 Structure
The rest of this work is structured as follows. In the next section, we describe our methodology and the concept behind each attack. Next, Section 3 details our experimental results in various malware monitoring environments. Then, in Section 5, we first perform a root cause analysis of our findings and propose mitigation issues. Finally, we conclude the article, summarizing our contributions and proposing ideas for future work.
Environment Availability Type Attack
Bitdefender GravityZone EDR Proprietary EDR
Bitdefender GravityZone Sandbox Proprietary Malware sandbox
CAPEv2 Open source Malware sandbox
REDACTED REDACTED REDACTED
Kaspersky EDR Proprietary EDR
Microsoft Defender Proprietary EDR
Recorded Future Triage Proprietary/Cloud Malware sandbox
Velociraptor Open source Digital forensic & incident response
VirusTotal Proprietary/Cloud Multiscan aggregator
Wazuh Open source EDR
WatchGuard EDPR Proprietary EDPR
Hybrid-analysis Proprietary/Cloud Malware sandbox
Table 1: Overview of our results.
2 Concept and methodology
Figure 1: Overview of our methodology
Our goal is to stress the telemetry pipeline, not to disable sensors. Thus, the core concept of our approach is that some malware analysis mechanisms and anti-malware solutions, in order to properly record all the actions of malware or a process in general, follow greedy approaches, since one cannot estimate which actions would be performed, and there is no specific threshold. While anti-malware solutions may kill a process or even prevent its launch due to static features, the same does not apply to malware sandboxes, where users expect to log all the actions of a binary. However, allowing arbitrary writes can obviously backfire in various ways. Given that some filtering is already performed, we focused on the data representation.
Let us consider a trivial example of a malware that is executed in a malware sandbox. Its execution triggers some processes for which the malware sandbox would timestamp, keep track of the parent and children processes or even store their memory dumps. The natural way to represent this information is through a tree; however, if you need to share this information with others or store it in a database, this data structure might not be the most convenient way. This convenience may be easily achieved by using a JSON representation of the tree through nesting. Yet, arbitrary nesting is not natively supported by all libraries or databases.
The above example illustrates that a custom recursive spawning mechanism can reveal many flaws that use this mapping. To this end, we consider a properly crafted malware binary that repeatedly launches embedded copies of itself, incrementing a depth counter with each iteration until a predetermined maximum depth is reached. Once the maximum depth is achieved, a lightweight payload (in our tests, a PowerShell reverse shell) is executed to confirm end-to-end execution. This design enables controlled stress on process-tracking, event logging, and downstream serialization systems.
The recursion state is tracked via environment variables, so the depth persists across spawned processes. The implementation evolved from a minimal proof-of-concept in C/C++ to a hardened test binary incorporating multiple stages of XOR-based encryption; this multi-stage design reduces trivial static detection, allowing us to test EDRs and ensuring the dynamic behavior appears only at runtime.
Figure 1 summarizes the method’s workflow and the primary failure modes exercised. On the left are the targeted environments (sandboxes and endpoint agents) in which a proof-of-concept binary was executed. The central column represents the operational limits we triggered in our experiments: recursion limit refers to code-level or timeout protections that abort deep traversal; oversized output captures large JSON or BSON blobs that exceed serializer or database limits; database thresholds indicate document size and nesting caps or insert limits that cause errors; and speed limit symbolizes UI rate caps that produce agent queue overflows or unresponsive dashboards. The panel to the right shows the telemetry mechanisms we targeted (process and call-argument data structures and persistent storage), where failures manifest as truncated or missing process trees, corrupted or absent behavioral reports, and UI stalls. Together, the elements illustrate how a simple recursive load can propagate through the telemetry pipeline and produce denial-of-analysis conditions.
2.1 Threat model and assumptions
In the following paragraphs, we detail our threat model and assumptions in order to frame our attack and its potential impact better. Thus, we list all the parameters that are relevant for our attack, what is controlled and by whom, but also what we expect to achieve.
Execution context.
We assume that an adversary can store an unprivileged user-mode binary on either (i) an endpoint that is instrumented by sandboxing or EDR/DFIR tooling, or (ii) at a cloud solution, by uploading the binary through common user interfaces. The binary recursively spawns child processes up to a chosen depth and rate. Execution occurs under typical analysis conditions (e.g., detonation VMs, lab hosts, or enterprise-managed endpoints) where process creation is permitted and environment variables propagate to child processes.
Telemetry.
We assume that monitoring stacks capture process execution (e.g., via user-mode hooking and/or kernel/event providers), serialize these events into structured records (e.g., JSON/BSON or comparable encodings), and store them to a database-backed backend that exposes dashboards or APIs to analysts. Default configurations of such tools (e.g., document size caps, recursion/nesting limits, per-index write limits, dashboard rate caps, and analysis timeouts) are used as-is and not tuned to any specific workload. Also, the adversary cannot modify or disable sensors, hooks, or ETW/Sysmon providers; load kernel drivers; or gain elevated privileges, directly access or reconfigure serialization libraries, databases, or dashboards.
We assume the telemetry pipeline ingests data derived from untrusted program behavior (e.g., deep process trees, long command-line arguments, large argument arrays, and nested object graphs), and that serialization and storage components treat these data as valid inputs rather than as isolation boundaries.
Resource bounds.
We assume that recursive spawners use a maximum depth chosen to avoid gross host-level denial of service (e.g., kernel scheduler collapse). We do not rely on exhausting OS resources (e.g., CPU, RAM, PIDs) as the primary failure mechanism. This was proven to be true on all tested occasions.
Network and isolation.
According to our evaluation methodology, execution can occur in offline or network-restricted environments. A network is not required for the core effect of the presented methodology.
Non-assumptions.
The attack does not assume administrative privileges, kernel drivers, agent tampering, or vendor-specific vulnerabilities. It does not require disabling sensors, unhooking instrumentation, or modifying backend configurations.
Non-goals.
This work does not aim to achieve persistence, privilege escalation, data exfiltration, or evasion of dynamic behavior capture. We do not claim cryptographic bypasses, supply-chain compromise, or exploitation of memory corruption. Likewise, we did not introduce obfuscation mechanisms that would completely blind the payload, but only the necessary ones to prevent EDRs from blocking execution from static analysis. As such, we only focus on introducing operational failures in telemetry pipelines using valid but manipulated inputs generated from the data sources they are programmed to collect.
3 Experimental process and environment setup
In what follows, we detail our experimental process, discussing the setup, goals, metrics, environment, and our results.
Listing 1: Nested JSON for spawned processes.
1{
2 "processes": {
3 "id": "id1",
4 "timestamp": 1761675797,
5 "children": [
6 {
7 "id": "id2",
8 "timestamp": 1761675798,
9 "children": [
10 {
11 "id": "id3",
12 "timestamp": 1761675797,
13 "children": [
14 {
15 "id": "id4",
16 "timestamp": 1761675799,
17 "children": [...]
18 ]
19 }
20 ]
21 }
22 ]
23 }
24 ]
25 }
26}
3.1 Setup
As described in our methodology (section 2), the idea is that we create a binary which gradually removes encryption layers by spawning a decrypted version of itself. As such, in each execution, the binary keeps a counter of the spawned processes to determine when to stop. At that step, our binary would open a reverse shell to a host that we controlled. Thus, we intend to create a structure similar to 1, where nesting or associated storage would create storage and/or processing issues. Minor changes to the binaries were introduced to explore the potential problems as we evaluated each software solution.
The experiments followed a progressive strategy. We started with open-source sandboxing platforms to iterate quickly and observe internal failures. After validating the concept on those systems, we moved to open-source EDR and DFIR tools and locally deployed trial instances of commercial products. Finally, sanitized samples were submitted to public, cloud-based analysis services to validate behavior with online providers. All local testing was performed in isolated, offline laboratory networks using virtual machines, to prevent telemetry leakage and to ensure safe handling of test artifacts. Each sandbox or EDR instance was installed with default or documented configurations; where supported, we varied timeouts, process depth limits, memory dump options, and report compression settings to exercise different code paths. Tests were repeated multiple times to ensure reproducibility; observed failures were cross-validated across independent environments when possible. Logging and monitoring of host resource usage (CPU, memory, disk I/O) were performed for each run to correlate system load with observed telemetry failures. The isolation of the experimental environment from the internet was introduced to avoid the communication of telemetry to software vendors that would allow them to prevent the execution of our binaries by adding some static checks or patching their solution on the fly.
3.2 Goals and evaluation criteria
The primary objective of the experimental phase was to validate whether recursive process spawning could cause denial-of-analysis or otherwise impair the behavioral visibility of sandbox and endpoint detection systems. The experiments were designed to observe three main outcomes: (i) denial or corruption of analysis reports, (ii) partial or total loss of behavioral telemetry, and (iii) degradation of analysis performance or stability leading to timeouts or crashes.
Each local test environment, whether sandbox, EDR, or DFIR platform, was evaluated using identical input parameters, including recursion depth and number of spawned processes. The consistency of execution ensured that the differences in outcomes reflected genuine architectural or implementation differences between the platforms rather than sample variance. The evaluation was carried out on five main pillars:
  1. 1.
    Behavioral completeness: The degree to which each platform successfully captured and recorded all executed behaviors (process creation, PowerShell invocation, network connections). Missing or truncated telemetry indicated a partial failure.
  2. 2.
    Detection verdict: The final classification or confidence score assigned to the sample (e.g., clean, suspicious, malicious). A low confidence score despite active payload execution was considered an analysis failure.
  3. 3.
    System stability: Observable signs of engine instability, including crashes, unresponsive interfaces, incomplete reports, or unprocessed data queues.
  4. 4.
    Report integrity: Verification that the JSON, BSON, or database-backed output was correctly generated and rendered without corruption, truncation, or serialization errors.
  5. 5.
    Performance overhead: Comparative measurement of processing time, report generation delay, and resource utilization under recursive load compared to baseline runs with benign samples.
We considered a test successful if the sample prevented the behavioral engine from generating a complete or accurate report or caused instability in the telemetry or storage pipeline. Conversely, a platform was considered resilient if it fully recorded the recursive activity and correctly labeled the behavior as malicious without disruption.
3.3 Test environments and safety controls
Our methodology included testing across a variety of platforms to exercise different telemetry and persistence architectures:
  1. 1.
    Open-source sandboxes. <REDACTED> and CAPEv2 were used to evaluate behavior logging and process-tree handling. These platforms rely heavily on JSON-based reporting and were instrumental in identifying truncation, serialization, and database size limitations.
  2. 2.
    Open source EDRs and monitoring tools. Wazuh was used to analyse agent-based monitoring under high-volume recursive process creation (truncated alerts, Sysmon event loss, memory pressure). Velociraptor evaluated forensic and live-response visibility under recursive load.
  3. 3.
    Commercial trial and local deployments. Tools such as Bitdefender GravityZone, Kaspersky EDR, WatchGuard EDPR, and Comodo Valkyrie were deployed in offline virtual machines to assess detection resilience without external telemetry leakage.
  4. 4.
    Online sandboxes. After offline validation, sanitized samples were submitted to public services including VirusTotal, AnyRun, Hybrid-Analysis, Malwation Threat Zone, Recorded Future Triage, OPSWAT Filescan.io, and Comodo Valkyrie to verify behavior across cloud providers.
All test systems were instrumented to capture system events, agent logs, exported reports, and backend errors. Where available, raw JSON/BSON dumps, database error logs, and UI traces were retained for post-mortem analysis. Proof-of-concept artifacts and any reproduction materials were handled in accordance with institutional disclosure policies and stored on isolated lab infrastructure.
3.4 Payload obfuscation
To reduce the likelihood of early static detection from EDRs, the sample employs a multi-layer ‘‘onion’’ encryption scheme. Each runtime stage decrypts only the subsequent stage. This prevents the full payload or final command-and-control instruction from being visible during static inspection. The obfuscation is intentionally lightweight, sufficient to bypass basic static heuristics while keeping behavior reproducible across testbeds.
4 Experimental results
Due to variability in results across products, we detail the experimental results in the following paragraphs, categorizing them by attack success and product, and discussing the root causes for each product where appropriate.
4.1 Failed attacks on EDRs
Most EDRs we tested terminated the corresponding processes relatively early, whereas others completed the decryption and terminated the process once a connection to the attacking machine was established. Based on feedback from some vendors, they implemented this measure because their products detected an excessive number of spawned processes within a short time; therefore, they were flagged as fork bombs, and the corresponding processes were terminated. Clearly, EDRs do not have to record all interactions of a binary and wait for a final verdict. Once a binary triggers specific alerts, the EDR immediately kills the process, so our attack did not manage to succeed in most EDR solutions
4.2 Wazuh
A special case of EDRs is Wazuh. Wazuh’s pipeline is heavily dependent on JSON: Agents serialize endpoint events into JSON artifacts (e.g., ossec-alerts.json) and forward them for indexing and correlation. In our tests, adversially structured telemetry caused those JSON streams and alert files to balloon (we observed ossec-alerts.json and related logs exceeding 100-200 MB), flooding the reporting and correlation pipeline. The result was not an obvious crash, but a silent loss of visibility; critical malicious activity was not logged or surfaced, many alerts were delayed or dropped, and numerous Windows Error Reporting files indicated local instability (stack-overflow exceptions) on the monitored hosts.
From an analyst’s point of view, this is a denial-of-analysis. Wazuh continued running but failed to produce useful, actionable telemetry, creating dangerous blind spots and false confidence. Therefore, the finding highlights a systemic risk for JSON-based EDR solutions, which can be exploited to hide part of their actions.
4.3 Undetected malicious behavior from malware sandboxes
In the case of some malware sandboxes, e.g., Bitdefender GravityZone Sandbox, Threat Zone, while there is a verdict, the solution does not manage to detect what the malware was doing and report it to the analyst. In fact, in some cases, the solution might not have broken, e.g., Threat Zone, but the verdict was that the sample was benign.
4.4 Velociraptor
Velociraptor presented a noticeably different failure mode compared to the malware sandboxes. The problem we observed was not a silent database rejection or a library exception, but rather an operational breakdown in Velociraptor’s collection and rendering pipeline when asked to materialize very large process trees. Two related reproducible issues were identified during our testing.
First, the Generic.System.Pstree artifact attempts to return the entire process tree as a single JSON row. Under our recursive-load experiments, this row becomes extremely large and, during transmission or queuing, exhausts the client’s memory buffers. The client process is then terminated by the local OS, producing the following observable behavior: the flow aborts with the message:
Flow not known - maybe the client crashed?
and the result folder for the collection contains only a tiny (4 KB) compressed artifact, indicating severe truncation of the expected output.
The root cause was confirmed by Velociraptor’s engineering team; the artifact encoded the entire tree in a single JSON row. A patch was merged (PR # 4454, Sep 17, 2025) [^2] Patch reference: https://github.com/Velocidex/velociraptor/pull/4454 with the commit message:
Bugfix: Limit number of entries emitted into process_tracker_tree() . The Generic.System.Pstree artifact puts the entire tree into a single JSON row which can exceed buffer size in some case. This causes problems in deeply nested trees.
Second, even when the backend successfully receives the raw event stream, for example, when Windows.Events.TrackProcesses is configured to forward updates, the Velociraptor UI, and gRPC transport hit practical limits. In our tests, the raw sysmon-derived events did contain all spawned processes, but attempting to display or transmit that dataset to the GUI produced the following gRPC error
received message larger than max (8732115 vs. 4194304)
causing the interactive view to fail and the analyst to lose visibility. Therefore, although the events were captured, it was practically invisible to the operator, who had to perform low-level interactions to collect them.
4.5 CAPEv2
For CAPEv2, we used the commit 52e4b43 from 2025-05-17. CAPEv2 uses MongoDB for storage and relies heavily on JSON/BSON serialization in Python. The above introduces various constraints. For instance, the use of MongoDB implies two different quotas. First, when storing analysis reports in MongoDB, there is a 16 MB BSON document size limit. Similarly, MongoDB enforces a maximum nesting depth of 100 levels. Therefore, if reports exceed the 16 MB quota or the nesting exceeds 100, the database rejects the insert, causing report generation to fail, regardless of whether the corresponding collector worked correctly.
Setting aside the underlying database, CAPEv2 also relies on Python’s orjson library to serialize results into JSON. However, the latter also has its own limits regarding deeply nested or recursive structures. Therefore, this is another potential vulnerability condition, as properly crafted samples may cause the serializer to exceed Python’s recursion limit, leading to serialization failures and aborted report generation.
We believe that some of these limitations were known to the developers of CAPEv2, as there are some integrated mechanisms to mitigate these issues, e.g., by progressively pruning large parts of the report with the removal of child nodes or entire process subtrees. However, even this pruning is insufficient when behavior complexity surpasses thresholds, resulting in repeated failures to save the report.
Our tests resulted in a Denial-of-analysis in reporting/mongodb.py and reporting/jsondump.py in CAPEv2 that allows attackers who can submit samples to cause incomplete or missing behavioral analysis reports by generating deeply nested or oversized behavior data that trigger MongoDB BSON limits or orjson recursion errors when the sample executes in the sandbox. Note that many systems depend on CAPEv2 dynamic analysis, e.g., VirusTotal uses it as one of its sandboxes. This issue has been assigned CVE-2025-61301 https://nvd.nist.gov/vuln/detail/CVE-2025-61301.
From a system perspective, system logs show MongoDB OperationFailure Code 15 and ‘‘Recursion limit reached’’ errors during JSON serialization. Moreover, final reports often contain no behavioral data or show ‘‘failed reporting’’, falsely making the sample appear benign or inactive. Table 2 shows evidence collected across multiple analysis runs (report sizes, logged processes, and errors).
Timeout Custom Duration Logged BSON Logs Report.json Errors/ Behavior
(sec) Settings (sec) Processes (Files / Size) Size (KB) Warnings Status
60 enforce-timeout 413 24 25 / 126 KB 613 Behavior OK
60 252 38 39 / 200 KB 923 Behavior OK
60 full-proc-memory-dumps, import-reconstruction-proc-dumps, enforce-timeout, syscall 230 42 43 / 230 KB 1054 Behavior OK
100 full-proc-memory-dumps, import-reconstruction-proc-dumps, enforce-timeout, syscall 330 61 62 / 318 KB 1482 Behavior OK
100 277 81 82 / 493 KB 2032 Behavior OK
120 full-proc-memory-dumps, import-reconstruction-proc-dumps, enforce-timeout, syscall 340 72 73 / 383 KB 1793 Behavior OK
120 88 89 / 475 KB 2228 Behavior OK
180 full-proc-memory-dumps, import-reconstruction-proc-dumps, enforce-timeout, syscall 414 90 91 / 492 KB 2356 MongoDB Code 15 No behavior
180 323 88 89 / 475 KB 2228 Behavior OK
200 413 119 120 / 645 KB 3174 MongoDB Code 15 No behavior
200 full-proc-memory-dumps, import-reconstruction-proc-dumps, enforce-timeout 144 / 726 KB Recursion limit reached, MongoDB Code 15 Failed reporting
300 513 146 147 / 779 KB 3676 MongoDB Code 15 No behavior
300 full-proc-memory-dumps, import-reconstruction-proc-dumps, enforce-timeout, syscall 167 / 892 KB Recursion limit reached, MongoDB Code 15 Failed reporting
Table 2: CAPEv2 analysis results showing correlation between task complexity, report size, and failure conditions.
4.6 Triage
Similar to CAPEv2, a denial-of-analysis vulnerability was found in Triage [^3] https://tria.ge/, a widely used commercial malware sandbox. Given that the platform is proprietary and the vendor did not disclose further details, we provide our own speculations. Nevertheless, this issue has been assigned CVE-2025-61303 https://nvd.nist.gov/vuln/detail/CVE-2025-61303.
Using the same concept as in CAPEv2, we submitted a sample to Triage for assessment. The sample generates a very high process volume and deeply nested activity. Using, for instance, Windows 10 v2004 and Windows 10 LTSC 2021 images to test our sample, a denial-of-analysis condition is performed, allowing a submitted sample to exhaust process-tracking and logging resources by recursively spawning large numbers of child processes. The result is truncated or missing telemetry; the PowerShell execution and reverse-shell stages that execute on the host are not recorded or reported. In fact, the initial behavioral output reported a score of 1/10, which marks it actually as benign, despite the sample executing PowerShell reverse-shell stages and other malicious actions on the host.
Reproducible public detonations are available on Triage [^4] https://tria.ge/250822-qfystshk51, which we used to validate the behavior across the provider’s Windows images. Based on the Triage’s API, we know that the behaviour analysis report is a JSON, which, for the processes, uses a nested format. Therefore, we assume that the same process is followed when collecting the events, leading to a crash that truncates the output. We do not consider that the issue is in the presentation layer, as a score is fetched (1/10), and part of the process tree is reported.
4.7 <REDACTED>
We cannot disclose the vendor name, product build identifiers, or specific commit references at this time, because the vendor requested an additional 30 day remediation window to deploy multiple patches across their systems. The vendor acknowledged the issue and is actively working on patching; we have submitted the report to MITRE for CVE assignment and are awaiting a CVE identifier. Per our coordinated disclosure agreement, details will remain private until the remediation window elapses. Therefore, for the remainder of the subsection we refer to the affected platform simply as <REDACTED>.
For <REDACTED>, our tests resulted in a Denial-of-analysis in the behavior-logging and report-generation pipeline of <REDACTED>’s <REDACTED> deployments (including <REDACTED> commit <REDACTED> and <REDACTED> Sandbox X.X.X used by <REDACTED> and <REDACTED>) allows attackers who can submit samples for analysis to cause incomplete or missing behavioral reports by triggering unbounded recursive process creation that generates deeply nested or oversized behavior data. The excessive log volume can exceed the JSON or database serialization limits, leading to silent analysis failure or a frozen web interface.
We reproduced the issue on both public instances and locally cloned deployments. In the affected analyses, we regularly saw <REDACTED> and <REDACTED> grow to 45-50 MB, so the underlying collectors were clearly logging a lot of activity; however, the final behavioral report was empty. Attempting to open those reports often caused the web UI to stall or the browser to freeze, and in offline clones, the large report files and UI instability prevented access to otherwise completed analyses. Representative runs and measures artifacts are summarized in Table 3.
Timeout Processes Total Size <REDACTED> <REDACTED> <REDACTED> <REDACTED> <REDACTED> UI behaviour
30s 5,000 $\sim$2.9 MB 68 KB 2.6 MB 183 KB 422 B 120 KB No lag
60s 5,000 $\sim$7.7 MB 151 KB 7.1 MB 471 KB 580 B 304 KB No lag
100s 5,000 $\sim$20 MB 387 KB 18 MB 1.3 MB 1.4 KB 823 KB Lags; $\sim$10 s to load
120s 5,000 $\sim$22 MB 397 KB 20 MB 1.4 MB 2.2 KB 846 KB Lags; $\sim$15–20 s to load
180s 5,000 $\sim$34 MB 687 KB 31 MB 2.3 MB 1.7 KB 1.5 MB Browser tab unresponsive
240s 5,000 $\sim$35 MB 693 KB 32 MB 2.3 MB 2.2 KB 1.5 MB Unresponsive from beginning
300s 5,000 $\sim$46 MB 894 KB 40 MB 3.0 MB 2.2 KB 1.8 MB Unresponsive from beginning
400s 5,000 $\sim$51 MB 999 KB 46 MB 3.4 MB 3.4 KB 2.2 MB Tab crashed
Table 3: Representative runs and measured artifacts for <REDACTED>.
5 Discussion
5.1 Root cause analysis and mitigation measures
Based on our findings, several practical steps can reduce denial-of-analysis risk across monitoring platforms and the impact of TCAs.
Nesting every element in a single JSON object is suboptimal, as most languages and serializers impose depth and size constraints. Mapping processes to tree structures or graph databases can alleviate these issues; otherwise, large or deeply nested data should be split into chunks so that partial results are saved even when limits are exceeded, preventing full report failure. Telemetry should also be validated early, at the agent or collector boundary, against a strict schema and sanitized before ingestion. When information is truncated, a standardized marker should indicate alteration so analysts can identify and recover affected data. To preserve auditability, platforms should maintain an isolated, access-controlled forensic store that accepts raw artifacts when the main pipeline rejects them.
In the same vein, we highlight the crucial role of proper queue management and event collection. Sudden spikes in the monitored events and the corresponding load of recording the associated information can cause many problems, e.g., dropped events or collector crashes. Of special interest are short and temporary or recursive processes that poor queue management may miss, leading to, e.g., losing the relationship between child and parent process, making them appear as orphans. We believe that delaying the processing of new processes, especially in malware sandboxes, could throttle such spikes and allow the collector to properly manage the collected information and the high number of events generated by spawned processes. Similarly, they may aggregate repetitive events (e.g., compress process-spawn sequences into summarized records) and canonicalize equivalent objects to cut down volume before transmission.
The reporting mechanisms should also become more robust. For example, incremental rendering of dashboards and reports can prevent interface issues with oversized datasets and selectively provide truncated previews when needed. In most cases, not all information has to be displayed to the user at once; therefore, some information can be prioritized, hinting that further information exists, improving usability and stability. An assessment of how much information exists and how this can be properly loaded, given the memory limits, internal size, and recursion thresholds, can also facilitate this process. Thus, dashboards must render lazily, using, e.g., virtualized trees and pagination, and avoid attempts to materialize large hierarchies in the browser or server render threads. This can be facilitated by using streaming JSON/BSON parsers to avoid loading entire documents into memory, enabling partial consumption and checkpointing of long streams.
5.2 Limitations
While TCAs introduce a new set of denial-of-analysis conditions for antimalware solutions, we acknowledge several limitations of our work. First, although we tested multiple antimalware solutions, the results may not be generalizable to all commercial or proprietary solutions, as differences in architecture, telemetry pipelines, and database backends could yield varying outcomes. Evidently, serialization libraries, the underlying data structures, as well as the goals that each solution has, affect the applicability of our attacks. For instance, an EDR is expected to be less prone to such an attack, as a fast forking process follows an aggressive pattern that the host does not need to be exposed to, so killing the process after a specific and small threshold is more optimal.
Second, the observed failures were primarily specific to systems relying on Python, MongoDB, or JSON/BSON serialization, meaning that platforms using different languages, serialization formats, or database engines might display distinct thresholds or error behaviors. Of course, this is mostly relevant for the open-source solutions that we were able to locally test and monitor each step of the process. For commercial solutions, since they are a black box to us, we observed only the external behavior; internal implementation is unknown. Nevertheless, this may also mean that some relevant vulnerabilities exist, and we missed them either because we did not use the correct threshold or because the output or missed detection was a result of our attack.
Finally, while we propose some high-level mitigation strategies such as back-pressure mechanisms, incremental reporting, and robust serialization, a complete implementation and empirical evaluation of these defenses across diverse products, as well as the impact on other factors, e.g., efficiency and latency, are beyond the scope of this work.
6 Related work
Dynamic analysis of binaries with and without the use of automated sandboxing is central to malware analysis. Prior work has shown various ways in which adversaries try to defeat these systems. Evasion of malware sandboxes usually aims at the identification of the sandbox environment through various means, such as timing attacks against virtualization, registry, network, or CPU artifact attacks, and human interaction detection [4]. For example, SandPrint showed that sandbox implementations may leak stable fingerprints that malware can detect and react to [25]. Timing and resource channels remain practical. Embedding proof-of-work into malware exposes timing asymmetries between bare-metal and scaled cloud sandboxes, reducing the effectiveness of online analysis services [17].
6.1 Attacks on EDRs
While EDR systems offer far more comprehensive visibility than traditional antivirus solutions, they are not immune to evasion or abuse. Prior research and practitioner analyses have demonstrated multiple ways to bypass [10, 19, 9] or even weaponize them [6, 1].
One of the main directions in this research line targets the hooking mechanisms used by EDRs to intercept user-mode or kernel-mode APIs. Several works exploit unhooking or direct system-call invocation to execute unmonitored code paths, typically by calling native Windows functions from ntdll.dll without passing through instrumented stubs [2, 9, 14]. Because many EDRs rely on user-land hooks, both attackers and defenders operate at the same privilege level, creating inherent race conditions, which at the end favor the attackers. Recent work, such as Cymulate’s Blindside technique, employed hardware breakpoints to selectively allow the loading of ntdll.dll while blocking other libraries. [^5] https://cymulate.com/blog/blindside-a-new-technique-for-edr-evasion-with-hardware-breakpoints/
Another common evasion vector leverages legitimate binaries and scripts to execute malicious payloads, known as living-off-the-land (LotL) tactics. [^6] https://lolbas-project.github.io/ Attackers may also abuse DLL sideloading; placing a malicious library next to a trusted executable so that it loads attacker-controlled code [10]. [^7] https://github.com/reveng007/AQUARMOURY Control Panel (.cpl) files have likewise been repurposed to execute arbitrary payloads outside of standard inspection paths. [^8] https://github.com/Tylous/ScareCrow, https://github.com/rvrsh3ll/CPLResourceRunner.
The aforementioned methods have been widely used in the real world, as well as methods to terminate the EDR processes. Indeed, some attackers may resort to the use of vulnerable drivers to terminate the corresponding processes, as in the case of GhostEngine [^9] https://www.elastic.co/security-labs/invisible-miners-unveiling-ghostengine. Moreover, following the Malware-as-a-Service paradigm, cybercriminals sell tools like Бафомет EDR Killer [^10] https://www.eset.com/blog/en/business-topics/threat-landscape/stop-edr-killers/, which automate this process and disable EDRs. Likewise, more advanced attackers have been reported manipulating Windows Defender Application Control (WDAC) policies to disable EDRs or whitelist malicious binaries  [^11] https://beierle.win/2025-08-28-A-Nightmare-on-EDR-Street-WDACs-Revenge/.
More offensive strategies attempt to blind EDRs by blocking or corrupting their telemetry. For instance, tools such as EDRSilencer use Windows Filtering Platform (WFP) rules to cut off network traffic to EDRs, [^12] https://github.com/netero1010/EDRSilencer while EDR-Freeze [^13] https://www.zerosalarium.com/2025/09/EDR-Freeze-Puts-EDRs-Antivirus-Into-Coma.html exploits Windows Error Reporting to suspend the monitoring processes.
6.2 Attacks on malware sandboxes
Contrary to EDRs, malware sandboxes allow malware to execute, as their goal is to actually observe what the malware does when it is executed. Therefore, malware tries to hide as much information as possible. Therefore, malware needs to identify the environment in which it operates and determine whether it will continue its operation. For instance, since the malware sandbox is a virtualized environment, malware tries to determine whether they are being executed in a virtual machine. Side information, e.g., names of devices or running processes, hardware identifiers, screen resolution, available memory, or even available threads and processors can easily give away the execution in a virtualized environment [5, 21, 12]. Similarly, a malware may try to find drivers, registry keys, or even folder names associated with virtual machines. Going a step further, malware may try to find elements of human interaction to determine whether it is executed in a malware sandbox. For instance, the uptime, the lack of keyboard and mouse interaction, the absence of browser history, or recent file activity can be used to infer the environment. Therefore, many malware would perform these checks before execution of the actual payload. Moreover, this creates a very fragile and sensitive environment where different configurations can provide radically different reports.
Given the resource and time constraints, some malware may opt to delay their payload execution [11]. This forced malware sandboxes to manipulate time to trigger their execution faster. As a result of this arms race, malware tries to measure the timing of specific functions to determine whether the elapsed time is aligned with the expected processing time. Evidently, time plays an important role in the dynamic analysis of malware [13].
For more on evasive malware, the interested reader may refer to [8].
6.3 Denial-of-Service attacks on security solutions
Petsios et al. presented SlowFuzz, a method to automatically synthesize inputs that trigger worst-case for several well-known algorithms, causing severe behavioral issues in widely deployed libraries [18]. For what is relevant in this work, many various zip parsers used in antivirus software whose decompression was significantly delayed. Other related work detected and repaired DoS-prone regex patterns, often used by cybersecurity solutions to detect strings or bytes associated with malicious activity (e.g., ReDoSHunter, RegexScalpel) and exposed vulnerabilities in non-backtracking matchers [15, 24, 16]. In content-inspection pipelines, adversarial inputs that induce super-linear or exponential matching can exhaust CPU budgets, delay processing, and cause packet or alert drops.
6.4 Positioning
Our attacks are conceptually related in spirit to classic structured-data and resource-exhaustion techniques such as the ‘billion laughs’ XML entity expansion and fork bombs. Although the ‘billion laughs’ attack was initially targeted at the XML parsers [23], similar attacks can be launched against other parsers, e.g., YAML [20] and SAML [22]. Likewise, fork bomb attacks target resource limits by rapidly spawning processes [3]. Thus, we sit between these two ideas. Like billion-laughs, we exploit pathological data structures, and like fork bombs, we exploit process creation to exhaust downstream resources. However, we try to achieve completely different goals on different systems and architectures. In this regard, we consider our work to be closely related to the recent work of Filic et al. [7], in which the authors exploit implementations of well-known probabilistic data structures in various solutions.
More precisely, our focus falls in the Denial-of-Service space against antimalware and malware sandboxes, but differs from current approaches in that we target weaknesses in data structures and telemetry-handling paths of monitoring systems. Instead of hiding or escaping containment, we weaponize the components used in data handling to induce denial-of-service or analysis failures. Not only do we effectively demonstrate that these types of structures are widely used in many industry anti-malware solutions, but, to the best of our knowledge, we also show that this attack surface is underexplored and follows a distinct line of research, including evasion, regex-based availability, and other attacks.
7 Conclusions
This work introduces TCAs, a novel method for inducing denial-of-analysis in modern anti-malware and telemetry monitoring systems by exploiting weaknesses in their data-handling and storage pipelines. We demonstrated that behavioral telemetry, not just sensors or hooks, can be an attack surface. Unlike traditional techniques that focus on hiding malicious behavior or evading detection, we directly target the internal telemetry infrastructure, demonstrating that unbound collection, serialization, and storage of monitoring data can itself be weaponized to disrupt analysis. Indeed, across 12 widely deployed products and cloud services, seven exhibited some form of denial-of-analysis or behavioral invisibility. Thus, two vendors assigned CVE identifiers (CVE-2025-61301 and CVE-2025-61303), others patched or reconfigured their software, and vendors acknowledged the issue during coordinated disclosure.
As malware analysis and detection increasingly rely on automated pipelines, availability becomes a security property. The recursive workloads in our experiments often pushed JSON and BSON serialization to their limits, suggesting a need for more thorough stress testing of deeply nested and complex data structures to identify safe thresholds and improve the robustness of serialization pipelines. Moreover, our results highlight the importance of viewing telemetry pipelines as potential attack surfaces and hardening them accordingly. Thus, in future work, we plan to explore fuzzing of telemetry structures and graph-based storage for behavioral data. Moreover, since Python and MongoDB were at the core of most observed denial-of-analysis conditions, it is essential to test other programming languages, serialization libraries, and database engines to check whether similar weaknesses exist.
Acknowledgements
This work was partially supported by the European Commission under the Horizon Europe Programme, as part of the projects SafeHorizon (Grant Agreement no. 101168562) and LAZARUS (Grant Agreement no. 101070303).
The content of this article does not reflect the official opinion of the European Union. Responsibility for the information and views expressed therein lies entirely with the authors.